OpenStudy (anonymous):

The tallest player on JEfferson High's basketball team is 6 1/2 feet. How many inches taller than 50 inches is he?

OpenStudy (anonymous):

2.33 feel

OpenStudy (anonymous):

78=x-50

OpenStudy (anonymous):

feet*

OpenStudy (anonymous):

How did you solve it? If it is not too much to ask. Thank you

OpenStudy (anonymous):

convert 6 1/2 ft to inches, 78 inches 78=x-50, that is how to solve